1) when the temperature-time data graph was reviewed, it was found that the error had been made determining DeltaT. Instead of 4.76C, Delta T was actually 4.70C. Based on this change only, calculate the correct deltaH neutralization for the reaction oh NH3 and acetic acid. 2) calculate percent error if for correct Delta H (neutzn) if Delta T 4.76C had been used? <-----thew second question is what I want
Answer #1

Answer 8 - Roge no-10 (1) let, Volume of C-13CCOH = 500 ml volume of NH3 = 500ml concentration of CH2C00H = LOM or mor ConcenPage No - (2) Given, Дт: •16. QE MXOYOT = 100 of x4.184 I x4,216d = 1991.58uJ At Neuta = 1991-584 T 0.05 MOI = 398.31.685 mo